# Low-rank orthogonalization for large-scale matrix optimization with applications to foundation model training

### v1

@@ 4.3 LLaMA pretraining
In this subsection, we consider pretraining LLaMA [50], a transformer-based language model with a more refined architecture than GPT-2. We experiment with LLaMA models of sizes 60M, 135M, 350M, and 1B on the same three datasets tested in the Muon GitHub repository [28]: FineWeb10B, FineWeb100B, and FineWebEdu10B.
Figure 7: Singular value distributions of the Q, K, and V matrices: across layers at iteration 300 (top row) and over training iterations at layer 16 (bottom row) during LLaMA pretraining.
@@
We apply low-rank Muon, Muon, AdamW, and SGD with momentum (SGDM) for pretraining LLaMA. Similar to Section 4.2, we implement two versions of low-rank Muon with rank parameters $100$ and $200$, and we use AdamW to train the embedding and head layers for Muon and low-rank Muon. We initialize all methods with the same weights from pretrained LLaMA models, and terminate all methods after one training epoch, consisting of 5000 iterations. We compare all methods using validation perplexity. The algorithmic parameters are carefully tuned to suit each method well in terms of computational performance. More details about the experimental setups including algorithm and LLaMA model parameters are deferred to Appendix B.
We visualize the top 100 singular values of the momentum updates $\{M^{k}\}$ associated with the Q, K, and V matrices for a 350M LLaMA model trained by Muon on the FineWeb100B dataset in Figure 7. The first and second rows of the figure display the top 100 singular values across different training iterations and neural network layers, respectively. This visualization partly supports the low-rank nature of the momentum updates.
@@
We present a comparison of the validation perplexity and computational time in Table 2. From this table, we observe that our low-rank Muon consistently achieves better validation perplexity compared to Muon for the majority of model sizes and tested datasets, while vastly improving upon the validation perplexity achieved by AdamW and SGDM. We also observe that the computational time of our low-rank Muon improves upon that of vanilla Muon, though all Muon-type methods remain slower than SGDM and AdamW. These observations demonstrate that our low-rank orthogonalization produces more generalizable LLaMA models across all tested model sizes and datasets. Moreover, it reduces computational time compared to the Newton-Schulz iterations for orthogonalization, although the improvement is less significant than the gains discussed in Sections 3.1 and 4.1. This is because most of the computational time is dominated by the forward and backward passes of the neural networks.

### v2

@@ 4.2 LLaMA Pretraining
In this subsection, we consider pretraining LLaMA (Touvron et al., 2023), a transformer-based language model with a more refined architecture than GPT-2. We experiment with LLaMA models of sizes 60M, 135M, 350M, and 1B on the same three datasets tested in the Muon GitHub repository (Jordan et al., ): FineWeb10B, FineWeb100B, and FineWebEdu10B.
@@
GaLore 94.90 86.99 86.49 44.71 $2.49\text{e3}$ $5.27\text{e3}$ $1.20\text{e4}$ $2.77\text{e5}$
SOAP 50.51 40.17 35.76 24.14 $2.46\text{e3}$ $5.27\text{e3}$ $1.21\text{e4}$ $2.78\text{e5}$
@@
GaLore 58.97 45.14 37.05 35.56 $2.65\text{e3}$ $5.50\text{e3}$ $1.09\text{e4}$ $2.69\text{e5}$
SOAP 55.08 52.28 47.52 43.89 $2.64\text{e3}$ $5.50\text{e3}$ $1.07\text{e4}$ $2.70\text{e5}$
@@
GaLore 64.04 68.13 45.13 43.56 $2.57\text{e3}$ $5.27\text{e3}$ $1.20\text{e4}$ $2.76\text{e5}$
SOAP 40.32 28.10 27.79 24.54 $2.60\text{e3}$ $5.29\text{e3}$ $1.20\text{e4}$ $2.78\text{e5}$
@@
We apply low-rank Muon, Muon, AdamW, GaLore, SOAP and SGD with momentum (SGDM) for pretraining LLaMA. Similar to Section 4.1, we implement two versions of low-rank Muon with rank parameters $100$ and $200$, and we use AdamW to train the embedding and head layers for both Muon and low-rank Muon. We initialize all methods with the same weights from pretrained LLaMA models, and terminate all methods after one training epoch, consisting of 5000 iterations. We compare all methods using validation perplexity. The algorithmic parameters are carefully tuned to suit each method well in terms of computational performance. More details about the experimental setups, including algorithm and LLaMA model parameters, are provided in supplementary materials.
@@
We present a comparison of validation perplexity and computational time in Table 2. From this table, we observe that our low-rank Muon consistently achieves better validation perplexity compared to Muon, GaLore, and SOAP for the majority of model sizes and tested datasets, while vastly improving upon the validation perplexity achieved by AdamW and SGDM. We also observe that the computational time of our low-rank Muon improves upon that of vanilla Muon, though all Muon-type methods remain slower than SGDM and AdamW. These observations demonstrate that our low-rank orthogonalization produces more generalizable LLaMA models across all tested model sizes and datasets. Moreover, it reduces computational time compared to the Newton-Schulz iterations for orthogonalization, although the improvement is less significant than the gains discussed in Section 3.1, as most of the computational time is dominated by the forward and backward passes of the neural networks.